March 29, 2017, Production & Analysis Working Group

From GlueXWiki
Jump to: navigation, search

Meeting Time and Place

The meeting will be on Wednesday March 29, 2017 at 2:00 pm EDT. For those people at Jefferson Lab, the meeting will be in room F326-327.

Meeting Connections

  1. To join via a Web Browser, go to the page [1]
  2. To join via Polycom room system go to the IP Address: ( and enter the meeting ID: 115815824.
  3. To join via phone, use one of the following numbers and the Conference ID: 115815824.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. More information on connecting to bluejeans is available.

Reconstruction & Analysis Studies Status


  1. Announcements
  2. Data Production
  3. Analysis Focus
  4. This Week's Studies
  5. Any other studies
  6. Upcoming Study Schedule

This Week's Topics

  • Talks: Summary/overviews only. Detailed discussions should be held in the appropriate working groups.
  • Talks should be limited to 10 + 5 minutes.

Data Production

  1. Monitoring Update --- Thomas Britton
  2. Calibration Update --- Sean Dobbs
  3. Preparing for the first Spring 2017 production launch
  4. Processing Update --- Alex Austregesilo

Analysis Focus: Total Cross Sections

  1. J/Ψ, φ --- Lubomir
  2. η, ρ, ω, φ --- ω cross section for several amorphous runs with tagger coincidence--Simon
  3. φ --- Thomas
  4. Any others?

Studies: Beamline & Triggering

Updates on experiment, simulation, and the comparison between the two.

  1. Flux --- Justin
  2. Beam energy --- Beamline Group
  3. Polarization (TPOL & lineshape) --- Beamline Group
  4. Beam Asymmetries --- Alex Austregesilo
  5. Trigger emulation --- Alex Somov
  6. Triggering efficiency --- Alex Somov

Data Production Milestones

  1. Fri. April 7 - Finalize calibrations/software for low-rate running. Run final tests.
  2. Fri. April 14 - Start production for low-rate running.
  3. Fri. April 21 - Finalize calibrations/software for high-rate running. Run final tests.
  4. Fri. April 28 - Start production for high-rate running.

Upcoming Study Schedule

  • Updates on experiment, simulation, and the comparison between the two.
  • Summary/overviews only. Detailed discussions should be held in the appropriate working groups.
  • All talks should be limited to 10 + 5 minutes.
  • Next Week: Hit Efficiencies

(i.e. Is the detector working?)

  1. CDC Hit Efficiencies --- Naomi Jarvis
  2. FDC Hit Efficiencies --- Alex Austregesilo
  3. BCAL Hit Efficiencies --- Elton Smith
  4. FCAL Hit Efficiencies --- Jon Zarling
  • Week after next: Alignment & Track/Shower Efficiencies

(i.e. Is the reconstruction working?) Updates on experiment, simulation, and the comparison between the two.

  1. Drift chamber alignment --- Mike Staib
  2. Tracking Efficiencies --- Paul Mattione
  3. BCAL Shower Efficiencies --- Elton Smith
  4. FCAL Shower Efficiencies --- Jon Zarling
  • The following week: Other
  1. Track / Hit Matching: BCAL, FCAL, TOF, SC --- Paul Mattione
  2. Efficiencies: TOF, SC
  3. Means & Resolutions (time, energy, dE/dx): Tracking, BCAL, FCAL, SC, TOF
  4. Uncertainties: PID (BCAL, FCAL, TOF, dE/dx), Kinfit (BCAL, FCAL, tracking)
  5. Channel/Analysis Studies: Branching ratios, cross sections, SDMEs, beam asymmetries
  6. Other reconstruction/analysis issues


Data Production

  1. Sean reported that calibrations are progressing. The main ongoing activities are tagger calibrations, CDC gains, calorimeter non-linearities, and drift chamber alignment. The other high priorities are to check the stability of calibrations and other outputs from the latest monitoring launch.
  2. Paul reported the results of a first look at creating an additional reduced set of REST files that can be more easily shipped off-site. Using one low-rate run, if events with at least one track that are matched to a tagger photon in the microscope, the REST size was reduced by a factor of ~4.5. However, this will result in a small loss of signal, as it assumes that the correct RF bunch can be chosen, even though it may be biased by junk/accidental tracks.
  3. Alex A. reported on the latest (ver10) monitoring launch. Only the low-rate runs were processed. Some tagger calibration problems remain, Sean reported that these are almost resolved.
  4. Mark I. reported that simulation launch 1.2.1 has completed. It is for Spring 2016, 6 billion events, no EM background.

Analysis Focus: Total Cross Sections

  1. Simon followed up his ω cross section sfor several amorphous runs by adding in new results with tagger coincidence. The values are generally lower and have larger bin-to-bin variations. This could be related to a problem in tagger/PS matching, since the tagger efficiency cancels out in the cross section calculation when using the tagged flux. Simon has several things to look at.
  2. Thomas has seen some reasonable initial results for φ cross sections, but is working out some apparent inconsistencies before presenting them.

Studies: Beamline & Triggering

  1. Justin has made some preliminary looks at the fluxes from the monitoring launch, but did not have results to show.
  2. Sean has added the latest beam energies into the CCDB. Richard Jones will follow up with Alexandre Deur to see if there will be any more updates before the reconstruction launch.
  3. Nathan showed preliminary results from the TPOL analysis. Generally, the coherent peak position was stable, and after adding terms to take the photon beam being potentially off-center into account, he obtained consistents results for the average polarization across the run period using two different methods. The polarization between differetn diamond orientations sppears to differ significantly, more studies are planned.
  4. Alex Somov reported on the status of the trigger emulation. He has begun analyzing data takin in the last run period for consistency checks. His preliminary results are that the trigger efficiency for pi+pi- + p events is ~95%. He is working on implementing this calculation into the simulation.
  5. Alex A. reported on the results for rho reconstruction from the last monitoring launch. He i able to reconstruct several runs and extract polarizations, although his preliminary results are slightly lower than those from the TPOL analysis.